OpenStudy (anonymous):

The original price of a digital keyboard was $119.50. The keyboard was on sale for $109.94. What was the discount rate?

OpenStudy (anonymous):

9.56

OpenStudy (anonymous):

119.50 - 109.94 = 9.56 9.56/119.50 = 0.0802501 Round that to 0.081 0.081 * 100 = 8.1 The discount rate was 8.1%

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

yah thx!